OpenStudy (anonymous):

g(x) = -x^2+1 Determine the equation of g^-1(x) algebraically and state its domain and range.

OpenStudy (anonymous):

\[g(x)=-x^2+1\] Determine the equation of \[g^-1(x)\] algebraically and state its domain and range.

OpenStudy (anonymous):

To do this, replace y with x y = -x^2 + 1 x = -y^2 + 1 x - 1 = -y^2 1 - x = y^2 y = sqrt(1-x) so the inverse of g(x) is sqrt(1-x)